Let theta=310 degrees. Below we see a unit circle and a circle with radius 3 along with a radius representing theta. Using your calculator, what are the coordinates of both points? Thanks for the help!
Answer by Theo(13342)

about thre quickest and easiest way to do this is as follows:

hypotenuse of right triangle formed is the radius of the circle = 3.

the angle is formed from the origin of the graph to the surface of the circle.

use the following formula to find x and y.

cos(310) = x / 3

sin(310) = y / 3

3 is the hypotenuse.
cos is adjacent divided by hypotenuse.
sin is opposite divided by hypotenuse.

since you're in the fourth quadrant, sin is negative and cosine is positive.

solve for x and y in both those equations to get:

cos(310) = x/3 results in x = 3 * cos(310).

sin(310) = y/3 results in y = 3 * sin(310).

solve for x and y using your calculator to get:

x = 1.928362829

y = -2.298133329

your coordinate point on the circle will be (1.928, -2.298)

your coordinate point on the origin of the graph is (0,0).

the following graph shows this visually.

$$$